# The Surprising Effectiveness of Test-Time Training for Abstract Reasoning

### v1

@@ X BARC Ours 5353%
Li et al. (2024) found that program synthesis and fully neural predictors for ARC are highly complementary, even when trained on the same tasks. Their end-to-end neural model can only solve 42.2% of the tasks solved by the program synthesis model. However, we find that when equipped with our TTT pipeline, BARC’s fine-tuned fully neural model solves 73.5% of the tasks that are solved by the program synthesis model. This suggests that our TTT pipeline significantly improves the neural model’s ability to learn systematic reasoning patterns similar to those captured by program synthesis models.
@@ 7 Conclusion
In this work, we conduct an investigation of test-time training and demonstrate that it can significantly improve LM performance on the popular ARC dataset. We find that learning task-specific LoRA adapters and generating augmented test-time datasets using geometric transformations are crucial. We also develop an augmented inference pipeline that uses invertible transformations to generate multiple predictions and then uses self-consistency to select the best candidates. Our overall pipeline applies multiple test-time computation methods, with each component contributing positively. This suggests that not only can test-time compute improve LM performance, but different test-time methods can also complement one another. Our TTT pipeline, combined with an existing method (BARC), achieves state-of-the-art results on the ARC public set and performs comparably to an average human. Our findings suggest that test-time methods could play a pivotal role in advancing the next generation of LMs.
@@
Evaluation Framework
[…]
Experimental Reproducibility
Given the computational requirements of our experiments, this preprint reports results without comprehensive standard error analysis. Our preliminary observations indicate minimal variance across runs, and we plan to include detailed statistical analysis in the final version.
@@
Even though the base Llama-3 perform extremely poorly on the public validation set, the public availability of the dataset on various platforms (GitHub, Kaggle) introduces the possibility that these models may have encountered these examples during pre-training.
@@ Acknowledgments
We thank Aniruddha Nrusimha for helpful discussions on parameter efficient training, and Jyo Pari for feedback on early drafts of this paper. This work was supported by the National Science Foundation under grants IIS-2212310, IIS-2238240, and CCF-2217064.
@@ Appendix A ARC Dataset
We present the tasks in development set, format and evaliation for the ARC dataset (available in this https link.).
@@ A.1 Data Format
We use numpy’s array printing format for all the experiments as shown in Fig.˜8.
@@ A.2 List of 80 Tasks Used For Development
We use following (Table˜2) tasks validation tasks for our development.
Table 2: Selected development tasks and their hardness level based on LeGris et al. [2024b].
@@ A.3 Evaluation
We follow the competition rules and in any of the two pass@2 predictions of the system is correct, we consider that test as correct. In the reported task level accuracies, we did not give partial points if all tests are not solved, except the final table Table˜1.
@@
Appendix B TTT
We present transformation used in TTT and the training details.
B.1 Transformations
@@
We provide the augmentations used in TTT in the Section˜3, please refer to our code base for their implementations. After application of these augmentation, we additionaly shuffle colors and shuffle training examples. Note that these transformations are applied to all input and output grids.
@@
Table 3: We provide the augmentations use in our TTT procedure with their function signature and description.
@@
Reflect(0, reverse=True) Flips a grid horizontally and prepend to the left of the original grid
Reflect(1, reverse=True) Flips a grid vertically and prepend to the above of the original grid
Reflect(0, reverse=False) Flips a grid horizontally and append to the right of the original grid
Reflect(1, reverse=False) Flips a grid vertical and append to the left of the original grid
RandomTranslateXY() Shifts a grid randomly both in horizontal and vertical directions. The maximum shift size is 4
Transpose() Reflect a grid on diagonal
IncreaseResolution(2) Upscale the grid by interleaving elements in both horizontal and vertical directions
IncreaseHeight(2) Upscale the grid by interleaving elements in vertical direction
IncreaseWidth(2) Upscale the grid by interleaving elements in horizontal direction
@@
B.2 Training Setup & Hyperparameters
Table 4: TTT hyper-parameters
@@ B.2 Training Setup & Hyperparameters
optimizer AdamW [Loshchilov and Hutter, 2018]
We use torchtune[torchtune maintainers and contributors, 2024] library to train LoRA adapters on Llama-3 family of models. We apply LoRA training to query and value projection weights of the self-attention layer, to the MLP weights and to the output projection layer (was only available for Llama-3 8B in torchtune). We present hyper-parameters of this training in Table˜4.
Appendix C Inference
We resort to vLLM [Kwon et al., 2023] library for prediction as it provides fast kernels and batched inference for our models and LoRA inference. We just use greed decoding as we did not see improvements with temperature sampling in our early experiments. We use 90, 180 degree rotations, horizontal, vertical, and diagonal (transpose) flips as our invertible transformations.
@@ Appendix D LM Data Generation
We described three approaches in Section˜5 to use LM, we generated 6426 task generators by few-shot prompting GPT-4 and GPT-4o models [OpenAI, 2023, Hurst et al., 2024].
@@ D.1 Getting Descriptions for Tasks
This procedure is shown in Fig.˜10. We initially described 10 training tasks with the hierarchical-style shown in Fig.˜6. Then, for other training tasks tasks, we obtained less quality crowd-worker annotations from LARC [Acquaviva et al., 2022] project. By using our high-quality seed annotations and their LARC version, we 10-shot prompt and LM to produce high quality annotations for the other training tasks.
@@ D.2 Few-shot Prompting Details
Figure 10: Generating quality seed descriptions: We use few-shot prompting to generate descriptions for a given task, using 10 manually created seed descriptions along with crowd-worker annotations from Acquaviva et al. [2022] as few-shot examples. For a given new task, we similarly provide the LM with examples and crowd-worker annotations (available only for training tasks).
Appendix E Fine-tuning
In each dataset described in Section˜5, we generated approximately $600000$ ARC tasks from the available task generator functions (e.g. training tasks, ReARC tasks, and LM generated tasks) by repeatedly picking 2-7 examples from the generated input outputs, and also applying geometric transformations if used.
E.1 Fine-tuning Transformations
We use all the transformations given in Table˜3, and some additional transformations given in Table˜5. In the fine-tuning case, different from TTT, we apply augmentations to only inputs, only outputs or both.
Augmentation Name Description
DropoutOutput Randomly deletes some patches of the output grids.
DropoutInput Randomly deletes some patches of the input grids
E.2 Fine-tuning Hyper-parameters
We perform full fine-tuning on LLama-3 family models by using torchtune library. We train each model up to 16000 steps. We use 2xNVIDIA A100 GPU for 1B models, 4xNVIDIA A100 GPU for 3B and 8B models. We present hyper-parameters in Table˜6.
Table 6: Fine-tuning hyper-parameters

### v2

@@ X BARC Ours $53$%
Li et al. (2025) found that program synthesis and fully neural predictors for ARC are highly complementary. Their end-to-end neural model can only solve $42.2\%$ of the tasks solved by the program synthesis model. However, we find that when equipped with our TTT pipeline, BARC’s fine-tuned fully neural model solves $73.5\%$ of the tasks that are solved by the program synthesis model. This suggests that our TTT pipeline significantly improves the neural model’s ability to learn systematic reasoning patterns similar to those captured by program synthesis models.
Semi-private evaluation
ARC-AGI challenge provides a hidden “semi-private dataset” and performs external tests for submissions. We submitted our ensemble solution to the official ARC-AGI semi-private evaluation and observed $47.5$% accuracy. This decline may be attributed to more significant distribution shifts in the semi-private evaluation dataset. If the semi-private evaluation results become publicly available in future, we will provide a detailed analysis of these performance differences.
5 BIG-Bench Hard
5.1 Background
BIG-Bench Hard (BBH; Srivastava et al., 2023; Suzgun et al., 2023) is a benchmark comprising $27$ challenging tasks across $23$ task types, designed to evaluate large language models on reasoning, compositionality, and generalization. Unlike ARC, BBH features a broader natural language structure and lacks a shared input format, making it unsuitable for invertible transformations. However, this broader scope offers a valuable testbed for evaluating TTT’s effectiveness in a more generalized setting. Despite the absence of invertible transformations—previously used in ARC to expand the TTT dataset and enhance inference—TTT still significantly improves performance on BBH.
5.2 Experimental Details
Model architecture & optimization
We use Llama 3.1 (8B; Llama Team, 2024). For each task $d$, we train a separate set of LoRA parameters at test-time, with a LoRA rank of $64$ over $40$ random shuffles of the demonstration pairs to produce leave-one-out in-context tasks. More hyperparameter details are given in Section F.1.
On BIG-Bench Hard, our base language model is able to achieve non-trivial scores out-of-the-box. Consequently, we do not perform any initial fine-tuning on synthetic tasks outside of BBH like we do for ARC. Furthermore, since models achieve nonzero performance in a zero-shot setting, we provide the zero-shot results and analyze how TTT and ICL improve upon them.
Evaluation
For the $27$ tasks in BBH, we consider the $10$-shot setting, where we select $10$ random pairs from each task’s dataset to be demonstration pairs and evaluate on the remaining data. Each of the $27$ tasks is analogous to a single ARC task, consisting of $10$ labeled examples as demonstration pairs given at test-time. We report average results over five random seeds, where each seed specifies which $10$ examples form the demonstration subset. For more control over the evaluation process with test-time training, we write our own evaluation function, which is available in our codebase (for more details, see Section F.1). The number of evaluation examples for each task is then $240$ for all tasks except three: Causal Judgment, Penguins in a Table, and Snarks, which have $177$, $136$, and $168$ evaluation examples respectively. Note that the large number of evaluation samples for each task compared to ARC means we can do a task-specific analysis to analyze which types of tasks benefit the most from TTT (Section 5.4). Unlike with ARC, we do not have a collection of invertible transformations to run augmented inference. Instead, we use greedy decoding. Further hyperparameter details and evaluation details are given in Section F.2.
5.3 Impact of TTT Design
Figure 8: Overall BIG-Bench Hard Results. TTT outperforms standard in-context learning by $7.3$ absolute percentage points, from $50.5\%$ to $57.8\%$. Our performance improvement over direct input-output data shows that using in-context leave-one-out tasks is crucial. Not taking demonstration loss or taking loss on inputs results in a performance decrease. Unlike with ARC, using a shared adapter across all tasks improves performance.
In this section, we evaluate our method and its ablations, primarily comparing the zero-shot baseline, ICL, and TTT. No Example Permutation updates the model on a single in-context prompt instead of multiple shuffled versions. Direct I/O treats each input-output pair as separate training instances. Shared TTT uses a single adapter across tasks instead of task-specific adapters. No Demonstration Loss removes the loss applied to demonstration outputs. Loss on Inputs and Outputs extends the loss calculation to both inputs and outputs. These ablations are as detailed in Section 3. As these results are averages over $5$ runs, the standard errors of the mean for each method are given in Section F.1, averaging $0.4\%$.
The results in Figure 8 show that TTT achieves an overall accuracy of $57.8$%, outperforming standard ICL ($50.5$%) and Direct I/O learning ($51.5$%). This demonstrates that TTT’s capabilities extend beyond ARC to more diverse and complex reasoning tasks, proving its effectiveness in a broader range of natural language problem-solving scenarios.
We observe that TTT without example permutations—performing multiple gradient steps on a single in-context prompt before inference—reduces accuracy to a still-impressive $55.7\%$. Computing the loss only on the test output lowers accuracy to $54.4\%$, while applying it to both inputs and outputs achieves $55.9\%.$
Shared adapter
Unlike on ARC, using a shared adapter improves performance on BBH, indicating that tasks in BBH do not confound each other during training. On the ARC dataset, each puzzle has the same input format, so distinguishing among multiple tasks is difficult, and we may have conflicting gradients with a single adapter. In BBH, however, distinguishing tasks is trivial (the instructions differ in plain text), and many tasks are mutually helpful. For instance, updating on Logical Deduction Five Objects also aids Logical Deduction Three Objects, without hurting Word Sorting. Although this is no longer test-time training on distinct tasks presented individually at test time, it can be interpreted as TTT on the entire dataset presented collectively at test time.
5.4 Task-Specific Analysis
Our task-specific results show that performance improvements from TTT are highly task-dependent. Among the $27$ tasks in BBH, TTT results in a performance decline of at least $2\%$ compared to ICL in only $2$ tasks. In contrast, $12$ tasks show an improvement of at least $2\%$, with $9$ of these showing improvements of at least $5\%$. The four tasks with the most significant performance boost from TTT over ICL or zero-shot and the task with the most significant performance decrease are shown in Figure 9. These tasks in order of TTT’s improvement over ICL are Dyck Languages (parentheses matching), Ruin Names (humorous name modifications), Movie Recommendation (choosing similar films), Hyperbaton (adjective ordering), and Boolean Expression (evaluating a boolean expression). Detailed results for every task are given in Section F.2.
Figure 9: BIG-Bench Hard results for tasks with the largest TTT-ICL score differences. The four tasks on the left show the most significant improvements with TTT over ICL, while the task on the right has the lowest TTT score relative to ICL. Full task-specific results are given in Section F.2.
We hypothesize that improvements from TTT may be driven by tasks involving distribution shifts and structured patterns. For example, tasks like Dyck Languages and Hyperbaton follow clear grammatical or programmatic rules, which could align well with TTT’s ability to adapt to latent structural regularities during test-time.
Conversely, tasks requiring explicit step-by-step computation show limited gains with TTT. For instance, Boolean Expressions declined from $85.7\%$ to $80.4\%$ under TTT. This task’s algorithmic nature—dependent on sequential reasoning rather than pattern-based transduction—and its likely pre-training exposure suggest TTT’s updates may not resolve its specific demands. While these particular observations align with our hypothesis, the reason certain tasks benefit more from TTT remains an open question.
6 Related Work
Test-time training
The idea of updating model parameters at test-time using instance-specific data traces back to early work on local learning (Bottou & Vapnik, 1992). More recently, Sun et al. (2020) propose a simple test-time self-supervision scheme to adapt an image classifier when facing distribution shifts. In language modeling, Hardt & Sun (2024) fine-tune on retrieved neighbors at test-time for notable gains, while Hübotter et al. (2025) optimize retrieval via active data selection.
ARC challenge
Abstraction and Reasoning Corpus (ARC; Chollet, 2019; Chollet et al., 2025) is a collection of extremely challenging few-shot visual reasoning problems. Most approaches to ARC fall into two main categories: program synthesis and fully neural. Program synthesis approaches (Butt et al., 2024; Wang et al., 2024; Li et al., 2025; Greenblatt, 2024) first try to find the transformation function $f$, and then apply it to the test example. Fully neural approaches (Veldkamp et al., 2023; Bober-Irizar & Banerjee, 2024) try to directly predict the output $y^{\textrm{test}}$, only implicitly modeling $f$. In this work, we use a fully neural approach, using an LM to predict the test outputs. Recent work has explored hybrid methods, leveraging inference scaling and deep learning-guided program synthesis (Greenblatt, 2024; Li et al., 2025). Similarly, we find that integrating our neural model with program synthesis improves performance.
@@ 7 Conclusion
We conduct an investigation of test-time training and demonstrate that it can significantly improve LM performance on abstract reasoning and few-shot learning tasks, namely the Abstraction and Reasoning Corpus (ARC) and BIG-Bench Hard (BBH). Our key contributions include a robust TTT framework with leave-one-out in-context task construction, the optimization setup, and the inference strategy after TTT. Our results reveal the potential of TTT to tackle novel reasoning tasks, suggesting significant promise for test-time methods in advancing the next generation of LMs.
@@
Optimization bias
In development of ARC, we used a set of 80 tasks for validation/ablation experiments. Standard hyper-parameters (learning rate, epochs) were optimized using this set, which might have introduced some bias.
@@
While the base Llama-3 performs poorly on the public validation set of ARC, the public availability of the dataset introduces the possibility that these models may have seen these examples during pre-training. Similarly, while the base model achieves reasonable performance on BBH, its public availability raises similar concerns.
@@ Acknowledgments
We sincerely thank the BARC team (Li et al., 2025) for their support and collaboration in ensembling our method with theirs, resulting in an official joint submission to the ARC public set. We thank Aniruddha Nrusimha for helpful discussions on parameter efficient training. This work was supported by MIT–IBM Watson AI Lab, and by the National Science Foundation under grants IIS-2212310, IIS-2238240, and CCF-2217064. This work also benefited from many conversations during the Simons Institute Program on Language Models and Transformers.
@@ Appendix A ARC Dataset
We present the tasks in the development set, the data format and evaluation details for the ARC dataset (available at this https link.).
@@ A.1 Data Format
We use numpy’s array printing format for all experiments as shown in Figure 10.
@@ A.2 List of 80 Tasks Used For Development
We use the following (Table 2) tasks validation tasks for our development.
Table 2: Selected development tasks and their hardness level based on (LeGris et al., 2024).
@@ A.3 Evaluation
We follow the competition rules that if any of the two pass@2 predictions of the system is correct, we consider that test correct. In the reported task-level accuracies, we did not give partial points if all tests are not solved, except the final table Table 1.
@@
Appendix B Fine-Tuning Before TTT
Figure 11: LLM based synthetic tasks generation: Given some seed task descriptions and task generator functions in Python, we generate more generator functions to produce novel tasks. We use three different approaches: (1) few-shot prompting with only generators, (2) few-shot prompting with generators and task descriptions, (3) two-stage approach: first generate free form descriptions, then condition on them to generate more generators (shown in Figure 12).
B.1 Preparing Fine-tuning Data
(Hodel, 2024) provides domain-specific language (DSL), ReARC, as well as the transformation $f_{i}$ that solves the task-$i$, and the data generation function $g_{i}$ that are implemented in this DSL for each training task in the $\mathcal{D_{\textrm{ARC}}^{\textrm{train}}}$ dataset. These functions enable sampling of new input-output pairs that maintains the same underlying transformation principle:
$d=(x,y)\sim\textrm{eval}(g_{i})$ (1)
where $d$ represents a newly generated input-output pair that can be solved using the same transformation function $f_{i}$ as the original task-$i$ [^4] We can verify the generated examples by asserting $f_{i}(x)=y$..
(a) Using Existing Generators
The generator functions $g$ in ReARC already provide an effective data augmentation tool by producing different instantiations of same tasks. We generate extra samples from these training tasks by running the code many times and randomly splitting these new examples ($d\sim\textrm{eval}(g_{i})$) to a set of train and test examples. These augmented examples are already provided with their DSL release.
(b) Few-shot Prompting an LLM
Additionally, we used several approaches to generate novel tasks using an LM (in our case, an ensemble of GPT4 and GPT4-o).
$g^{\prime}\sim\textrm{LM}(g_{1},g_{2},\dots,g_{m})$ (2)
where $g^{\prime}$ is a new generator function and ${g_{1},\dots,g_{m}}$ are existing generator functions (shown in Figure 11). We sample different $m$ examples by uniformly from existing training set. We repeat this process multiple times to get a good amount of tasks.
where $s_{i}$ represents the description of task $i$.
To get the task descriptions, we manually created seed descriptions for $10$ training tasks. These seed descriptions were then used to generate descriptions for the training and validation tasks through few-shot prompting. To increase diversity of tasks, we use task descriptions with hierarchical fields (category, summary, and description). The process of getting these descriptions is provided in Section D.1.
Instead of jointly generating task descriptions and function generations, we additionally deployed a two-stage approach (Figure 12 ) described as following:
$s^{\prime}$ $\sim\textrm{LM}(s_{1},s_{2},\dots s_{m})$ (4)
This approach first generates a task description $s^{\prime}$ and then conditions the generator creation on both existing task pairs and the new description. In total we collected 6426 generators with these LLM based approaches. We provide qualitative samples from these LM generated tasks in Figure 16.
(c) Geometric Transformations
  • Input grids only: $(x,y)\rightarrow(t(x),y)$
  • Output grids only: $(x,y)\rightarrow(x,t(y))$
  • Both input and output: $(x,y)\rightarrow(t(x),t(y))$
We use all the transformations given in Section C.1, and some additional transformations given in Table 3. In the fine-tuning case, different from TTT, we apply augmentations to only inputs, only outputs or both. These transformations are applied randomly to variants of tasks with $30\%$ of the time.
Augmentation Name Description
DropoutOutput Randomly deletes some patches of the output grids.
DropoutInput Randomly deletes some patches of the input grids
B.2 ARC Initial Fine-tuning Hyperparameters
We perform full fine-tuning on LLama-3 family models by using the torchtune library. We train each model up to 16000 steps. We use 2xNVIDIA A100 GPU for 1B models, 4xNVIDIA A100 GPU for 3B and 8B models. We present hyperparameters in Table 4.
Table 4: ARC Initial Fine-tuning Hyperparameters
Hyperparameter Search Space
learning rate 2.5e-5
epochs 2
batch size 32
optimizer AdamW (Loshchilov & Hutter, 2018)
scheduler Cosine LR Schedule with 2k warmup
B.3 Results
We perform full fine-tuning 1B, 3B Llama 3.2 instruction-tuned, and 8B Llama 3 instruction-tuned using augmented data. The format and training objective is same as the ones described for TTT in 3. Hyperparameter details are given in Section C.2. We do the following ablations for augmented data:
  1. 1.
  2. 2.
    All: We use all methods described in Section B.1, including ReARC, rule-based augmentation, and LM generation.
  3. 3.
    No-Geom: We remove geometric transformations from all tasks.
  4. 4.
Figure 13: Left: Accuracy when fine-tuning with different data sources. While all fine-tuned models perform similarly, their performance after TTT shows considerable variance. As expected, removing geometric transformations from the fine-tuning data reduces performance compared to the model trained on the full dataset. Surprisingly, excluding LM-generated data from fine-tuning actually outperforms the model trained on all data. Right: Performance results across different model sizes. As expected, performance of the base fine-tuned model improves with increasing model size, aligning with current scaling law trends. However, the scaling behavior after TTT is less clear. For instance, the final performance of the 1B and 3B models is identical after TTT. Full discussion in Section B.3.
We show results using different model sizes in Figure 13. Increasing the model size consistently improves FT performance, with the 8B model achieving the highest accuracy of $36\%$. We also observe that TTT effectively closes the performance gap for smaller models, with the 1B and 3B models achieving similar accuracy after TTT.
Appendix C TTT Transformations for ARC
We present the transformations used in TTT and the training details.
Figure 14: TTT dataset generation for a test task (Section 3.1): We start by creating leave-one-out tasks from the given training examples of the task. These tasks are then augmented through rule-based transformations to obtain the full TTT dataset. Finally, we train task-specific LoRA adapters on top of the base FT model.
C.1 Transformations
@@
We provide the augmentations used in TTT in Table 5, please refer to our code base for their implementations. After applying these augmentations, we additionally shuffle colors and shuffle training examples. Note that these transformations are applied to all input and output grids. The procedure for generating the dataset for TTT is shown in Figure 14.
@@
Table 5: We provide the augmentations used in our TTT procedure with their function signature and description.
@@
Reflect(0, reverse=True) Flips a grid horizontally and prepends to the left of the original grid
Reflect(1, reverse=True) Flips a grid vertically and prepends to above the original grid
Reflect(0, reverse=False) Flips a grid horizontally and appends to the right of the original grid
Reflect(1, reverse=False) Flips a grid vertically and appends to below the original grid
RandomTranslateXY() Shifts a grid randomly in the horizontal and vertical directions. The maximum shift size is 4
Transpose() Reflects a grid on diagonal
IncreaseResolution(2) Upscales the grid by interleaving elements in both horizontal and vertical directions
IncreaseHeight(2) Upscales the grid by interleaving elements in vertical direction
IncreaseWidth(2) Upscales the grid by interleaving elements in horizontal direction
@@
C.2 Training Setup & Hyperparameters
We use the torchtune(torchtune Maintainers & Contributors, 2024) library to train LoRA adapters on Llama-3 family of models. We apply LoRA training to query and value projection weights of the self-attention layer, to the MLP weights and to the output projection layer (was only available for Llama-3 8B in torchtune). We present hyperparameters of this training in Table 6. We also found that using quantized LoRA adapters (Dettmers et al., 2023) instead of standard (full-precision) LoRA leads to only a small drop in performance ($29\to 26$ tasks solved with the 1B-parameter model), making it a viable option in memory-constrained settings.
We resort to the vLLM (Kwon et al., 2023) library for prediction as it provides fast kernels and batched inference for our models and LoRA inference. We just use greed decoding as we did not see improvements with temperature sampling in our early experiments. We use 90, 180 degree rotations, horizontal, vertical, and diagonal (transpose) flips as our invertible transformations.
With that, the whole TTT and inference process takes approximately $12$ hours for $100$ randomly sampled validation tasks when using an NVIDIA A100 GPU.
Table 6: ARC TTT Hyperparameters. We find learning rate of 5e-5 the best for 1B and 3B models, and 1e-4 the best for 8B models.
@@ C.2 Training Setup & Hyperparameters
optimizer AdamW (Loshchilov & Hutter, 2018)
@@ Appendix D LM Data Generation
We described three approaches in Appendix B to use LM, we generated 6426 task generators by few-shot prompting GPT-4 and GPT-4o models (OpenAI, 2024; Hurst et al., 2024).
@@ D.1 Getting Descriptions for Tasks
This procedure is shown in Figure 15. We initially described 10 training tasks with the hierarchical-style shown in Figure 11. Then, for other training tasks tasks, we obtained less quality crowd-worker annotations from LARC (Acquaviva et al., 2022) project. By using our high-quality seed annotations and their LARC version, we $10$-shot prompt and LM to produce high quality annotations for the other training tasks.
@@ D.2 Few-shot Prompting Details
Figure 15: Generating quality seed descriptions: We use few-shot prompting to generate descriptions for a given task, using 10 manually created seed descriptions along with crowd-worker annotations from Acquaviva et al. (2022) as few-shot examples. For a given new task, we similarly provide the LM with examples and crowd-worker annotations (available only for training tasks).
Appendix E Augmented Inference Pipeline
E.1 Augmented Inference
For a given task with training examples ${(x_{k},y_{k})}_{k=1}^{K}$ and test input $x_{\textrm{test}}$, we use invertible geometric transformations to produce equivalent transformed versions of the task, as shown in Figure 5. Let $\mathcal{T}$ be some set set of invertible geometric transformations (e.g., rotations and reflections). For each transformation $t\in\mathcal{T}$, we apply $t$ to all training demonstrations and the test input and run our model with these transformed inputs. We then apply the inverse transformation to obtain the final prediction for that transformation.
$y_{t}$ $=t^{-1}(\tilde{y})$ (7)
We further augment our predictions by permuting the order of training examples. For each transformation $g$, we sample $n=2$ different permutations of the demonstration sequence, resulting in $n\cdot|\mathcal{T}|$ total predictions per task. This is to mitigate any bias in the model’s processing of the demonstration sequence. (Bober-Irizar & Banerjee, 2024) also find transpose and rotation is helpful to produce extra prediction candidates.
E.2 Ensembling Predictions (Voting Strategy)
  1. 1.
  2. 2.
Appendix F BIG-Bench Hard Details
F.1 Further Experimental Details
We write our own evaluation function for BIG-Bench Hard available in our codebase. We found that existing evaluation frameworks did not properly measure zero-shot performance due to insufficient answer-extraction parsing and answer-format prompting. We also wanted more control in splitting each individual task’s dataset into demonstration examples and evaluation sets. For all results, we average results over different selections of the $10$ few-shot examples with the following $5$ random seeds: $42,43,44,45,46$. The full TTT and inference process takes approximately 15 minutes on an NVIDIA A100 GPU.
The standard error of the mean for each method in Figure 8 over the $5$ seeds is given in Table 7.
Table 7: Standard Error of the Mean for each method in Figure 8.
Method Standard Error of the Mean
Zero-Shot 0.01
ICL 0.19
TTT 0.20
No Example Permutation 0.32
E2E 0.66
Shared TTT 0.72
No Demo Loss 0.69
Loss on Inputs and Outputs 0.35
We search over the following hyperparameters:
Table 8: BBH TTT Fine-tuning Hyperparameters